OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 3y = 10 4x + 5y = 14 2 1 -1 -2 help please:)

OpenStudy (anonymous):

y=2 x=1

OpenStudy (anonymous):

how did you get that??

OpenStudy (anonymous):

subtract the first equation from the second 4x+5y-4x-3y=14-10 2y=4 y=2 then you can get x from either one

OpenStudy (anonymous):

oh ok i see now i was trying something compleatly different thank you!!